Circular No. 8246

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ke IN MCG +06-22-9 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Cen-

ter for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003ke (cf. IAUC 8243), obtained by M. Calkins on Nov. 28.51 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-IIn supernova. The spectrum consists of a moderately blue continuum with strong hydrogen emission lines.

SUPERNOVAE 2003kg AND 2003kh A. V. Filippenko, R. Chornock, and R. J. Foley, University of Califor-

nia, Berkeley, report that inspection of CCD spectra (range 320–940 nm), obtained on Nov. 29 UT with the Keck-I 10-m telescope, reveals that SN 2003kg and SN 2003kh (IAUC 8245) are of type Ia, with spectral-feature ages (cf. Riess et al. 1997, A.J. 114, 722) of 0 ± 2 and 7 ± 2 days (respectively) past maximum brightness. In both objects, though especially in the latter, the Ti II/Si II 590-nm absorption feature is unusually strong relative to the Si II 615-nm line, characteristic of subluminous type-Ia supernovae. However, the Ti II 420-nm absorption trough often seen in such objects (e.g., SN 1991bg; cf. Filippenko et al. 1992, A.J. 104, 1543) is largely absent. The expansion velocities of the ejecta, measured from the Si II 615-nm absorption, are 13700 km/s for SN 2003kg and 10300 km/s for SN 2003kh. The redshift of the host galaxy of SN 2003kh is 0.0334, measured from narrow Hα emission.

Corrigendum. The declination for SN 2003kh on IAUC 8245 should read +58^{o}58′52′′.5.

SUPERNOVAE 2003ih AND 2003ki Filippenko and Chornock add that inspection of CCD spectra, ob-

tained as above on Nov. 29 UT, reveals that SN 2003ih (IAUC 8207) is of type Ib/Ic, 2 ± 1 months past maximum brightness; the near-infrared Ca II triplet emission is especially strong. SN 2003ki (IAUC 8245) is of type IIb, ∼ 3–4 weeks after the explosion; weak He I absorption lines are present in the spectrum, whose hydrogen Balmer lines exhibit P-Cyg profiles. The redshift of its host galaxy is 0.025, measured from narrow emission lines in an H II region.

2003 December 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green